OpenStudy (anonymous):

simplify/reduce the algebraic fraction: 6x-3y/2x-y

OpenStudy (anonymous):

\[\frac{ 6x-3y }{ 2x-y }=\frac{ 3(2x-y) }{ 2x-y }\] Factor top and bottom completely and cancel any factors which are common to the numerator and denominator.

OpenStudy (anonymous):

Thank you! can you help with a few more?

OpenStudy (anonymous):

perhaps...

OpenStudy (anonymous):

Add the Algebraic Fractions and Reduce to lowest terms: 2x+3/x^2-9+x/x-3

OpenStudy (anonymous):

in order to add rational expressions, you need the terms to have a common denominator. What is the common denominator? \[\frac{ 2x+3 }{ x^{2} -9}+\frac{ x }{ x+3 }\]

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

you need to know the factors of each denominator. Can you factor the first denominator?

OpenStudy (anonymous):

no I am completely lost...please explain

OpenStudy (anonymous):

\[x^{2}-9 = (x-3)(x+3)\] this is a difference of squares

OpenStudy (anonymous):

okay because half of 9 is 3 right

OpenStudy (anonymous):

well 9 divided by 2

OpenStudy (anonymous):

no... the square root of 9 is 3.

OpenStudy (anonymous):

do you know how to add and subtract fractions?

OpenStudy (anonymous):

yeah okay I understand

OpenStudy (anonymous):

do you go across?

OpenStudy (anonymous):

no... you need a common denominator. I think you need some help with factoring as well. Unfortunately, I have to leave right now. I'll be back in a while but I suggest you get some help with fractions and factoring. That's what you'll need to do these problems.

OpenStudy (anonymous):

okay thank you for your help